OpenStudy (anonymous):

A 1.7 × 10 3 kg elevator carries a maximum load of 876.7 kg. A constant frictional force of 4.3 × 10 3 N retards the elevator’s motion upward. What minimum power must the motor deliver to lift the fully loaded elevator at a constant speed 3.45 m/s? The acceleration due to gravity is 9.81 m/s 2 . Answer in units of kW

OpenStudy (anonymous):

First, let's define power, so we know what we have and what we will need to determine. \[P = F \cdot v\] We are given the velocity at which the elevator travels, now all we need to do it determine the forces involved. I spot only two, gravitation and friction. \[F = F_g + F_f = m \cdot g + F_f\]